# MultiHateLoc: Towards Temporal Localisation of Multimodal Hate Content in Online Videos

### v1

@@ 5. Conclusion
We introduced MultiHateLoc, the first framework for weakly supervised multimodal hate content localisation, addressing a task that has been largely overlooked in prior research. By combining dynamic modality selection, cross-modal attention, and a modality-aware MIL objective, MultiHateLoc effectively captures the sparse and asynchronous cues that characterise hateful content across video, audio, and text. Experiments on the HateMM and MultiHateClip datasets demonstrate that MultiHateLoc can produce accurate temporal localisation under video-level supervision. Comprehensive ablations verify the contribution of each component, and qualitative visualisations further show how the model identifies discriminative segments across modalities. Overall, this work opens a new research direction for fine-grained multimodal hateful video understanding under weak supervision. We hope this work inspires further exploration for safer and more responsible online media environments.

### v2

@@ 5. Conclusion
We introduced MultiHateLoc, the first framework for weakly supervised multimodal hate content localisation, addressing a task that has been largely overlooked in prior research. By combining dynamic modality selection, cross-modal attention, and a modality-aware MIL objective, MultiHateLoc effectively captures the sparse and asynchronous cues that characterise hateful content across video, audio, and text. Experiments demonstrate that MultiHateLoc can produce accurate temporal localisation under video-level supervision. Comprehensive ablations verify the contribution of each component, and qualitative visualisations further show how the model identifies discriminative segments across modalities. Overall, this work opens a new research direction for fine-grained multimodal hateful video understanding under weak supervision.